OpenStudy (anonymous):

solve ; (12x/(3x-3))+(1/3)=4/(x-1)

OpenStudy (anonymous):

12x/3x-3 + 1/3 Multiply the right hand term by x-1/x-1. 12x/(3x-3) + (x-1)/(3x-3)=4/(x-1) Now multiply the right hand side by 3/3 12x/(3x-3)+(x-1)/(3x-3)=12/(3x-3) (13x-1)/(3x-3) -= 12/(3x-3) Multiply both sides by 3x - 3 13x - 1 = 12 13x = 12 + 1 13x = 13 x = 1 If this helps, please choose good answer. Visit my website at: http://www.tutorsean.net for free math videos.